### v1

# arXiv:2603.01298v1[q-fin.PM] 01 Mar 2026

Abstract
This paper introduces methodologies for constructing an index composed of a risky asset and a risk-free asset that achieves a fixed target volatility. We propose a simple proportional-control-based approach for setting the index weights, and we demonstrate in simulation that this method is more effective at consistently achieving the target volatility than an open-loop approach. We additionally present a modification to our proportional control approach that reduces index drawdowns in simulation.
1 Introduction
We propose a method to form an index composed of two assets only—a risky asset and a risk-free asset—achieving a fixed pre-specified volatility. We work under two assumptions: first, the risky asset cannot be shorted and second, we have an upper limit on leverage.
Volatility‐targeting strategies trace back to the rule of scaling portfolio exposure inversely with a forecast of return variance, as formalized by [moreira2017]. Large empirical studies—e.g., [harvey2018]—confirm that such open-loop rules improve Sharpe ratios and reduce tail risk across many asset classes. However, they also highlight practical issues such as turnover, leverage spikes, and sensitivity to volatility estimation error.
Recent papers attempt to remedy these weaknesses. Conditional schemes restrict scaling to extreme volatility regimes [bongaerts2020], while Bayesian or exponentially-weighted moving average (EWMA) smoothing dampens weight oscillations [bernardi2022]. At the same time, [cederburg2020] document that out-of-sample benefits can vanish once implementation frictions are considered. In contrast, we cast volatility management as a simple proportional feedback-control problem that explicitly corrects tracking error and, in an extension, throttles leverage during drawdowns—all using transparent parameters.
2 Methodology
Our goal is to construct an index from a single risky asset and a single risk-free asset (for illustration, we use cash) which achieves a specified level of volatility $\sigma^{\text{tar}}>0$ without exceeding a limit $L>0$ on the weight allocated to the risky asset. We call $\sigma^{\text{tar}}$ the target volatility and $L$ the leverage limit. We denote by $t_{1},t_{2},\ldots$ the time sequence at which we observe the returns of the risky asset, realize interest on cash balances, and have the opportunity to rebalance the index. We define the trailing returns of the risky asset and risk-free asset at time $t_{k}$ as $r_{k},r^{\text{rf}}_{k}\in\mathbb{R}$ respectively.
Index weights and returns.
An index is a pair of weights $(w_{k},c_{k})$ at time $t_{k}$ such that $w_{k}\in\mathbb{R}$ is the weight of the risky asset and $c_{k}\in\mathbb{R}$ that of cash. The weights must satisfy the conditions
$w_{k}+c_{k}=1,\qquad 0\leq w_{k}\leq L.$
The weights represent the fraction of the index invested in the risky asset and cash, respectively. A negative cash weight corresponds to leverage, or borrowing cash to invest more in the risky asset, which we can take in a limited capacity. When taking leverage, we pay interest on the borrowed cash rather than accruing positive interest on the cash. Between two consecutive times $t_{k-1}$ and $t_{k}$, the index return $r^{\text{ind}}_{k}$ is given by
$r^{\text{ind}}_{k}=r_{k}w_{k-1}+r^{\text{rf}}_{k}c_{k-1}.$
Realized volatility.
The realized volatility $\hat{\sigma}_{k}$ of the risky asset at time $t_{k}$ is defined as the square root of an EWMA of the squared risky asset returns $r_{1},\ldots,r_{k}$,
$\left(\hat{\sigma}_{k}\right)^{2}=\frac{1-\beta}{1-\beta^{k}}\sum_{j=1}^{k}\beta^{k-j}\left(r_{j}\right)^{2},$ (1)
with a halflife $h>0$ and decay factor $\beta=\exp(-\log(2)/h)$ [johansson2023]. (This is a EWMA estimate of the second moment, but we use the fact that daily returns have means that are much smaller in magnitude than the returns.) Note that $\hat{\sigma}_{k}$ is known at the time that $w_{k}$ is decided.
Realized index volatility.
Analogously, we define the realized index volatility $\hat{\sigma}^{\text{ind}}_{k}$ at time $t_{k}$ as the square root of the EWMA of the squared index returns $r_{1}^{\text{ind}},\ldots,r_{k}^{\text{ind}}$,
$\left(\hat{\sigma}^{\text{ind}}_{k}\right)^{2}=\frac{1-\beta}{1-\beta^{k}}\sum_{j=1}^{k}\beta^{k-j}\left(r_{j}^{\text{ind}}\right)^{2}.$ (2)
Note that $\hat{\sigma}^{\text{ind}}_{k}$ is known at the time that $w_{k}$ is decided.
Tracking error.
We define our relative instantaneous volatility tracking error $e_{k}$ at time $t_{k}$ as
$e_{k}=\log(\hat{\sigma}^{\text{ind}}_{k}/\sigma^{\text{tar}}).$ (3)
The instantaneous tracking error $e_{k}$ is approximately the percentage difference between the realized index volatility $\hat{\sigma}^{\text{ind}}_{k}$ and the target volatility $\sigma^{\text{tar}}$. Over a period $t_{j},t_{j+1},\ldots,t_{k}$, the average absolute volatility tracking error is
$\frac{1}{k-j+1}\sum_{i=j}^{k}|e_{i}|.$
The volatility control problem.
The problem is to determine a policy that chooses $w_{k}$ and $c_{k}$ at time $t_{k}$ based on information known at time $t_{k}$ (such as $\hat{\sigma}_{k},\hat{\sigma}_{k}^{\text{ind}},e_{k}$) in such a way that the average absolute volatility tracking error is minimized.
2.1 Open-loop approach
In the simplest, ideal case, the returns of the risk-free asset are known constants and the returns of the risky asset are independently and identically distributed according to a normal distribution with mean $\mu\in\mathbb{R}$ and standard deviation $\sigma>0$ at each time $t_{1},\ldots$. In this case, the index returns at time $t_{k+1}$ are distributed according to a normal distribution with mean $w_{k}\mu+c_{k}r^{\text{rf}}_{k+1}$ and standard deviation $w_{k}\sigma$. If we assume that $|\mu|\ll\sigma$ and $r^{\text{rf}}_{k}\ll\sigma$, then the expected value of $(r_{k+1})^{2}$ is approximately $\sigma^{2}$ and the expected value of $(r_{k+1}^{\text{ind}})^{2}$ is approximately $w_{k}^{2}\sigma^{2}$. It follows by the linearity of expectation that the expected variance $(\hat{\sigma}_{k})^{2}$ of the risky asset is approximately $(\sigma)^{2}$.
In this case, a natural approach to choosing $w_{k}$ and $c_{k}$ such that the expected value of $(r_{k+1}^{\text{ind}})^{2}$ is equal to $\sigma^{\text{tar}}$ is to set
$w_{k}=\min\{\sigma^{\text{tar}}/\hat{\sigma}_{k},L\},\qquad c_{k}=1-w_{k}.$ (4)
When $\sigma^{\text{tar}}/\hat{\sigma}_{k}\leq L$, we set the risky asset weight $w_{k}$ so that the expected value of the index variance is equal to the target variance. When $\sigma^{\text{tar}}/\hat{\sigma}_{k}>L$, we set $w_{k}=L$, accepting that this is the best we can do given the leverage limit. We call this approach the “open-loop” approach because the weights $w_{k}$ are set without any regard to the previous index weights $w_{1},\ldots,w_{k-1}$.
This methodology is summarized in Algorithm 1.
Algorithm 1 Open-loop solution to index construction
1: Input: $\sigma^{\text{tar}},L$
2: for $k=1,2,\ldots$ do
3:   Observe risky asset volatility $\hat{\sigma}_{k}$
4:   Set risky asset weight $w_{k}=\text{min}\{\sigma^{\text{tar}}/\hat{\sigma}_{k},L\}$
5:   Set cash weight $c_{k}=1-w_{k}$
In practice, although this method may sometimes perform reasonably, it fares poorly when there are changes in the prevailing volatility of the risky asset. This is illustrated in §3.
2.2 Volatility control
We now modify the methodology from §2.1 to better react to shifts in the volatility of the risky asset. To achieve this, we add a control parameter $\kappa_{k}\in\mathbb{R}$ to our policy (4) and set $w_{k}$ and $c_{k}$ such that
$w_{k}=\text{min}\{\exp(\kappa_{k})\sigma^{\text{tar}}/\hat{\sigma}_{k},L\},\qquad c_{k}=1-w_{k}.$ (5)
Above, the term $\exp(\kappa_{k})$ scales the risk target and thus allows us to adjust for persistent overestimation or underestimation of the volatility of the risky asset. Scaling via $\exp(\kappa_{k})$ rather than $\kappa_{k}$ is a matter of convention. It is motivated by the observation that the volatility of a stock tends to vary on a logarithmic scale: for an asset with a mean volatility of $X\%$, periods of $2X\%$ and $\frac{1}{2}X\%$ volatility will be approximately equally likely.
It remains to discuss how to dynamically set the control parameter $\kappa_{k}$, and we describe a simple methodology, which depends on only four parameters. The four parameters are a proportional gain $g>0$, a control parameter lower bound $\kappa_{\text{min}}<0$, an upper bound $\kappa_{\text{max}}>0$, and a smoothing factor $\theta\in(0,1)$. After setting these parameters, we compute $\kappa_{k}$ via
$\kappa_{k}=(1-\theta)\text{clip}(-ge_{k};I)+\theta\kappa_{k-1}$ (6)
in which
$I=[\kappa_{\text{min}},\kappa_{\text{max}}],\qquad\text{clip}(t;[a,b])=\begin{cases}a,&t<a,\\ t,&a\leq t\leq b,\\ b&t>b.\end{cases}$ (7)
This controller is designed to adjust $\kappa_{k}$ based on the discrepancy between the realized volatility of the index $\hat{\sigma}^{\text{ind}}_{k}$ and the target volatility $\sigma^{\text{tar}}$. To simplify, assume $\theta=0$ (no smoothing) and that the volatility of the index has recently been lower than the target. Then the error $e_{k}$ will be negative, and the term $ge_{k}$ will be positive, meaning that we will target a higher volatility in the next iteration. Conversely, if the volatility of the index has recently been higher than the target, then $e_{k}$ will be positive, and $ge_{k}$ will be negative, meaning that we will target a lower volatility in the next iteration.
The gain parameter $g$ influences how reactive $\kappa_{k}$ is to discrepancies between the realized and target volatilities. The clipping imposes reasonable limits on $\kappa_{k}$ to prevent it from becoming too large or too small. The smoothing factor $\theta$ determines how quickly $\kappa_{k}$ can change; a larger value of $\theta$ means that $\kappa_{k}$ will change more slowly, while a smaller value implies $\kappa_{k}$ will change more quickly.
Since the volatility control methodology assumes our index has some history (for volatility estimation), we recommend either running or simulating the open-loop method for some time before switching to the volatility control methodology.
The volatility control methodology is summarized in Algorithm 2 and its performance is demonstrated in §3.
Algorithm 2 Volatility control for index weight selection
1: Input: $\sigma^{\text{tar}},L,g,\kappa_{\text{min}},\kappa_{\text{max}},\theta$
2: Set $\kappa_{0}=0$
3: for $t_{k}=t_{\text{start}},\ldots$ do
4:   Produce an index volatility estimate $\hat{\sigma}^{\text{ind}}_{k}$ from historical index returns
5:   Compute volatility control error as in (3)
6:   Compute volatility control parameter as in (6)
7:   Set risky asset weight $w_{k}=\text{min}\{\exp(\kappa_{k})\sigma^{\text{tar}}/\hat{\sigma}_{k},L\}$
8:   Set cash weight $c_{k}=1-w_{k}$
2.3 Leverage control
While producing an index realizing a fixed volatility is the primary objective, we would also like to mitigate drawdowns resulting from using leverage during market downturns. To achieve this, we introduce a methodology which sacrifices some volatility-tracking performance in order to attempt to reduce the leverage of the index during periods of negative returns. We achieve this by adding a control parameter $\kappa_{k}^{\text{lev}}\leq 0$ to the leverage limit $L$ in our policy (5), and set our weights such that
$w_{k}=\text{min}\{\exp(\kappa_{k})\sigma^{\text{tar}}/\hat{\sigma}_{k},\exp(\kappa_{k}^{\text{lev}})L\},\qquad c_{k}=1-w_{k}.$ (8)
We mandate $\kappa^{\text{lev}}_{\text{min}}\leq\kappa_{k}^{\text{lev}}\leq 0$, where $\kappa^{\text{lev}}_{\text{min}}<0$ is a fixed lower bound, so that we will only ever reduce the leverage limit $L$ or leave it as it is. Its purpose is to reduce the leverage limit $L$ when the index is trending downwards.
Let
$R_{k}^{\text{ind}}=\prod_{j=1}^{k}(1+r_{j}^{\text{ind}})$
be the cumulative return of the index at time $t_{k}$. From $R_{1}^{\text{ind}},\ldots,R_{k}^{\text{ind}}$, we compute EWMAs $\text{MA}_{k}^{\text{long}}$ and $\text{MA}_{k}^{\text{short}}$ of $R_{k}^{\text{ind}}$ with halflives $h^{\text{long}}>h^{\text{short}}>0$. We then set
$\kappa_{k}^{\text{lev}}=\text{clip}\left(-g_{\ell}\log\left({\text{MA}_{k}^{\text{long}}}/{\text{MA}_{k}^{\text{short}}}\right);[\kappa^{\text{lev}}_{\text{min}},0]\right).$ (9)
This has a simple interpretation: when the short-term moving average is smaller than the long-term moving average, the leverage limit is reduced. As earlier, the parameter $g_{\ell}>0$ is a gain that determines how quickly the leverage limit is reduced.
The leverage control methodology is summarized in Algorithm 3.
Algorithm 3 Volatility + Leverage control for index weight selection
1: Input: $\sigma^{\text{tar}},L,g,g_{\ell},\kappa_{\text{min}},\kappa_{\text{max}},\kappa_{\text{min}}^{\text{lev}},h^{\text{long}},h^{\text{short}},\theta$
2: Set $\kappa_{0}=0$
3: for $t_{k}=t_{\text{start}},\ldots$ do
4:   Produce an index volatility estimate $\hat{\sigma}^{\text{ind}}_{k}$ from historical index returns
5:   Compute volatility control error as in (3)
6:   Compute volatility control parameter as in (6)
7:   Compute leverage control parameter as in (9)
8:   Set risky asset weight $w_{k}=\text{min}\{\exp(\kappa_{k})\sigma^{\text{tar}}/\hat{\sigma}_{k},\exp(\kappa_{k}^{\text{lev}})L\}$
9:   Set cash weight $c_{k}=1-w_{k}$
In general, varying the leverage limit in order to reduce downside effects is a difficult problem since it is effectively equivalent to timing the market. Certainly, using the leverage controller does not guarantee smaller drawdowns or greater returns than using the volatility controller alone. In the example from §3, we see that the leverage controller works well, but this is not uniformly the case for all the risky assets we have tested. The leverage controller also seems more sensitive to the choice of the gain parameter $g_{\ell}$ than the volatility controller is to $g$, though we provide the parameter values that have worked well for us in our experiments in §3. We document the leverage controller because we feel it shows some promise, but we do not fully endorse it.
3 Results
In this section, we present results about the performance of Algorithms 1, 2, and 3.
Simulation details.
Our simulation spans the trading days from June 6th, 2000 to April 9th, 2025, inclusive. The risky asset is IVV, the iShares Core S&P 500 exchange traded fund (ETF), which provides broad exposure to large-cap U.S. equities. We use close-to-close returns for IVV and assume that returns are observed, new weights are computed, and the index is rebalanced instantaneously at market close. The return on cash is the daily Federal Funds Effective Rate. We compute the risky asset volatility estimate $\hat{\sigma}_{k}$ as the square root of an EWMA of the squared risky asset returns $r_{1},\ldots,r_{k}$ with a halflife of 126 trading days. We set the leverage limit to $L=1.5$ and the target volatility to $\sigma^{\text{tar}}=0.15/\sqrt{252}$, corresponding to an annualized volatility of $15\%$.
In Algorithm 2, we set $g=50$, $\kappa_{\text{min}}=-1$, $\kappa_{\text{max}}=1$, and $\theta=0.5$. In Algorithm 3 we additionally set $g_{\ell}=20$, $\kappa_{\text{min}}^{\text{lev}}=-2$, and otherwise use the same values for all the other parameters. We use a 126-trading-day EWMA of the cumulative index returns for $\text{MA}_{k}^{\text{long}}$ and a 42-trading-day EWMA for $\text{MA}_{k}^{\text{short}}$. For both Algorithms 2 and 3, we run the method for the first 10 trading days before engaging control. In both cases, we compute the index volatility estimate $\hat{\sigma}^{\text{ind}}_{k}$ as the square root of an EWMA of the squared index returns $r_{1}^{\text{ind}},\ldots,r_{k}^{\text{ind}}$ with a halflife of 126 trading days.
Discussion.
Figure 1 shows the cumulative returns of the IVV, the open-loop method, the volatility control method, and the volatility plus leverage control method. Figure 2 shows the running estimated volatility of IVV and that of the three methods over the same period. The Monte Carlo methodology described in Appendix A was used to compute 10th and 90th percentiles for the running volatility estimate of an asset with a true volatility of $15\%$.
Table 1 summarizes the performance of the four methods over the simulation period. The volatility tracking error is computed as the mean absolute error between the running volatility estimate of the method and the target volatility $\sigma^{\text{tar}}$, scaled by $\sqrt{252}$ to give an annualized value. Although IVV does not track a volatility target, we still provide a volatility tracking error for IVV for reference.
As shown in the figures and table, the leverage control method provides a slightly greater return and a slightly smaller maximum drawdown than the volatility control method. This performance does come at the cost of some tracking error with respect to the $15\%$ target. Both the leverage control and volatility control methods outperform the open-loop method in terms of return, volatility tracking error, and maximum drawdown.
Figure 1: Cumulative returns of four indices.
Figure 2: Running annualized estimated volatility of four indices. 15% target volatility is shown as a solid black line. A 90% confidence interval for the running volatility estimate of an asset with a true volatility of $15\%$ is shown as a shaded region.
IVV Open-Loop Volatility Control Leverage Control
Volatility Tracking Error $5.2\%$ $2.3\%$ $\mathbf{0.4\%}$ $0.6\%$
Annualized Return $9.0\%$ $7.4\%$ $9.0\%$ $\mathbf{9.7\%}$
Annualized Volatility $19.2\%$ $15.1\%$ $\mathbf{14.9\%}$ $14.7\%$
Sharpe Ratio $0.47$ $0.49$ $0.61$ $\mathbf{0.66}$
Maximum Drawdown $55.3\%$ $38.7\%$ $38.3\%$ $\mathbf{32.9\%}$
Table 1: Summary of performance metrics for four indices.
Acknowledgments
We would like to acknowledge Gabriel Maher, Rob Tibshirani, and Ludovic Breger for their work on an initial methodology we built upon. We would additionally like to thank Veronica Mai, Logan Bell, Alex Tzikas, Raphael Chinchilla, and Ron Kahn for helpful discussions and feedback.
Appendix A Evaluation
Suppose we produced an index that had a volatility exactly equal to $\sigma^{\text{tar}}$. Any empirical estimate of the volatility $\sigma^{\text{eval}}_{k}$ from realized returns will almost surely not equal $\sigma^{\text{tar}}$. It is thus important to understand how closely we can expect our volatility estimates to match the target volatility.
For this purpose, we suggest the following benchmarking methodology:
  1. 1.
    Independently sample $N$ (for some large integer $N$) synthetic return values $r_{1}^{\text{synth}},\ldots,r_{N}^{\text{synth}}$ from a normal distribution with mean 0 and standard deviation $\sigma^{\text{tar}}$.
  2. 2.
    Compute volatility estimates $\sigma^{\text{synth}}_{k}$ from the synthetic returns $r_{1}^{\text{synth}},\ldots,r_{k}^{\text{synth}}$ using the same method we used to compute $\sigma^{\text{eval}}_{k}$ from the index returns $r_{1}^{\text{ind}},\ldots,r_{k}^{\text{ind}}$.
  3. 3.
    Discard the first $n$ values of $\sigma^{\text{synth}}_{k}$ (where $n$ is a burn-in period for the volatility estimate).
  4. 4.
    Compute statistics of interest from the remaining values of $\sigma^{\text{synth}}_{n+1},\ldots,\sigma^{\text{synth}}_{N}$.
Specifically, because we do not have the ability to observe the true volatility of our index, this methodology allows us to quantify the uncertainty of volatility estimates and understand the fluctuations we can expect to see if the index is tracking the target volatility.
A.1 Example
Suppose we estimate the volatility of our index via a EWMA with halflife $h$. We define $\alpha=\exp(-\log(2)/h)$ to be the corresponding decay factor of the EWMA. If we had infinitely many synthetic returns values from a normal distribution with mean 0 and standard deviation $\sigma^{\text{tar}}$, then we can define the infinite-sample, halflife-$h$ EWMA volatility estimate
$\sigma^{\text{synth}}=\sqrt{(1-\alpha)\sum_{i=1}^{\infty}\alpha^{i-1}\left(r_{i}^{\text{synth}}\right)^{2}}.$
It is a result of probability that $\sigma^{\text{synth}}$ is approximately distributed according to a scaled $\chi$ distribution with $\frac{1-\alpha^{2}}{(1-\alpha)^{2}}$ degrees of freedom and a scale factor of $\sigma\frac{1-\alpha}{\sqrt{1-\alpha^{2}}}$. From this approximation, we can estimate the standard deviation of these of EWMA volatility estimates as a function of the halflife $h$. For $\sigma^{\text{tar}}=0.15/\sqrt{252}$, corresponding to an annualized volatility of $15\%$, we show the estimated standard deviation of $\sigma^{\text{synth}}$ as a function of $h$ in Figure 3.
Figure 3: Approximate standard deviation of the ideal EWMA volatility estimate $\sigma^{\text{synth}}$ as a function of halflife $h$ for $\sigma^{\text{tar}}=0.15/\sqrt{252}$.
Returning to the sampling-based methodology we proposed, we can verify that for reasonably large $N$, the synthetic volatility estimates $\sigma^{\text{synth}}_{n+1},\ldots,\sigma^{\text{synth}}_{N}$ will be distributed according to the distribution of $\sigma^{\text{synth}}$. For $\sigma^{\text{tar}}=0.15/\sqrt{252}$ and halflives $h=21,126$, we demonstrate this in Figure 4.
Figure 4: Distribution of EWMA volatility estimates $\sigma^{\text{synth}}$ from Monte Carlo simulation with $N=10000$ samples and the first $n=252$ discarded. The dashed lines show the approximate distribution of $\sigma^{\text{synth}}$ for halflives $h=21,126$.
As evidenced by Figures 3 and 4, EWMA volatility estimates of $\sigma^{\text{tar}}$ with longer halflives have significantly less variance. As such, we recommend using halflives of at least 63 when evaluating the performance of our index construction methods.